OpenStudy (anonymous):

(3-4 square root7)(3-4 square root7) I need help to approaching the the final answer.

OpenStudy (anonymous):

\[\left(3-4\sqrt{7}\times \right)\left(3-4\sqrt{7}\times \right)\] is this the question

OpenStudy (anonymous):